if the rate constant for a reaction tripled when the temperature rises from 3.00×10^2 k to 3.10 ×10^2 k.what is the activation energy of the reaction?
Answer #1

rate constant k2 = 3 k1

temperature T1 = 300 K

T2 = 310 K

ln (k2 / k1) = Ea / R [1 / T1 - 1/ T2]

ln (3) = Ea / 8.314 x 10^-3 [1/ 300 - 1/310]

Ea = 84.9 4 kJ /mol

activation energy = 85.0 kJ /mol